# Barriers for Learning in an Evolving World: Mathematical Understanding of Loss of Plasticity

### v1

@@ Abstract
Deep learning models excel in stationary data but struggle in non-stationary environments due to a phenomenon known as loss of plasticity (LoP), the degradation of their ability to learn in the future. This work presents a first-principles investigation of LoP in gradient-based learning. Grounded in dynamical systems theory, we formally define LoP by identifying stable manifolds in the parameter space that trap gradient trajectories. Our analysis reveals two primary mechanisms that create these traps: frozen units from activation saturation and cloned-unit manifolds from representational redundancy. Our framework uncovers a fundamental tension: properties that promote generalization in static settings, such as low-rank representations and simplicity biases, directly contribute to LoP in continual learning scenarios. We validate our theoretical analysis with numerical simulations and explore architectural choices or targeted perturbations as potential mitigation strategies.
@@ 1 Introduction
The extraordinary success of back-propagation in training deep neural networks often relies on two implicit assumptions. First, stationarity is assumed. This means that the data distribution encountered during training is similar to the distribution faced during deployment. As a result, post-training adaptation is minimal or absent. Second, a single random initialization of network parameters is the main source of diversity and exploration potential, a resource that is progressively consumed by optimization and not replenished. These assumptions falter when an artificial agent must operate and learn continuously within an environment characterized by changing dynamics or evolving task distributions. This scenario, commonly referred to as continual or lifelong learning, presents the stability-plasticity dilemma [Abraham and Robins, 2005, Chaudhry et al., 2018]. This dilemma demands that the system must be stable enough to retain previously acquired knowledge, yet plastic enough to effectively integrate new information.
Empirically, standard deep networks subjected to long sequences of tasks or slowly drifting data streams often exhibit a decline in their learning capability [Dohare et al., 2024, Berariu et al., 2021, Dohare et al., 2021, Nikishin et al., 2022, Lyle et al., 2023]. This phenomenon, termed loss of plasticity (LoP), is distinct from catastrophic forgetting [McCloskey and Cohen, 1989, Ratcliff, 1990, French, 1999], where new learning overwrites old knowledge. LoP specifically refers to the diminished ability to learn new information effectively over time. Common symptoms include exploding weight magnitudes [Nikishin et al., 2022], activation saturation, the emergence of “dead” ReLU units (whose upstream parameters cease to update) [Nair and Hinton, 2010, Sokar et al., 2023, Dohare et al., 2021, Lyle et al., 2022], a collapse in the effective rank of hidden layer representations indicating reduced feature diversity [Papyan et al., 2020, Huh et al., 2023, Kumar et al., 2020, Gulcehre et al., 2022], and redundancy or diminishing contributions from network components like attention heads or filters [Lyle et al., 2023]. Many of these issues have been highlighted in recent studies focusing on LoP [Dohare et al., 2023, Kumar et al., 2024, Ash and Adams, 2020].
Dohare et al. [2024] argue that such failures are intrinsically linked to the back-propagation algorithm itself. They posit that gradient descent optimized for transient single-task learning relies heavily on the initial random state for exploration, a resource that is consumed and not replenished during prolonged training. Their work demonstrates that standard deep learning methods can lose plasticity until they perform comparably to linear networks and suggests that maintaining plasticity requires mechanisms beyond pure gradient descent such as continually injecting diversity via methods like their proposed continual backpropagation.
Goal of this paper.
Motivated by observations about LoP, our paper revisits the dynamics of gradient descent and back-propagation through the lens of dynamical systems theory. We seek to answer the question:
What structural features inherent in gradient flow dynamics inevitably lead to LoP, and how might we design algorithms or architectures capable of perpetual adaptation?
The central theorem in this work is that the tendency of gradient-based optimization to favor low-rank or “simple” representations lies at the heart of plasticity loss. While properties like low effective rank and simplicity bias are often associated with improved generalization in the standard two-phase learning paradigm [Huh et al., 2023, Papyan et al., 2020, Zhang et al., 2017], we argue that these very properties become detrimental in continual learning settings. By reducing the effective dimensionality of the network’s feature space, they limit its capacity to adapt to novel information, thus contributing to the LoP observed by [Dohare et al., 2024] and others. While loss of plasticity (LoP) has been previously linked to hallmarks of low-rank representations in the literature, our work introduces a novel perspective and develops a formal framework that systematically unifies several previously disparate observations. In addition, our theory is the first to establish explicit connections between LoP and the geometries induced by learning dynamics and cloning, two active research domains that we argue hold considerable potential for advancing the study of LoP.
@@ 2 LoP Manifolds: Traps for Gradient Descent
In this section, we lay the groundwork for our analysis by defining Loss of Plasticity (LoP) within a dynamical systems framework and recalling the standard definitions for feed-forward neural networks and back-propagation. The stability of LoP manifolds, a crucial concept for understanding their persistence, will be discussed later in Sec.˜2. Let $\theta\in\Theta\subseteq\mathbb{R}^{p}$ represent the parameters of a neural network. We consider training on a stream of data $\{(x_{i},y_{i})\}_{i=1}^{N}$ using gradient descent or its stochastic variants. The objective is typically to minimize a loss $\sum_{i=1}^{N}\mathcal{L}(\hat{y}_{\theta}(x_{i}),y_{i}),$ which we can succinctly refer to as $\mathcal{L}(\theta)$. This allows us to define LoP based on the the trajectory of parameters $\theta(t)$ in the parameter space $\Theta$ as driven by the negative gradient in the loss landscape.
@@
Note that both LoP manifolds $\mathcal{M}_{F}$ and $\mathcal{M}_{C}$ are defined as linear LoP manifolds in the sense of Definition˜2.1. Formal proofs and further details are provided in Sec.˜A.2.
@@
It is important to note that the Duplicate Manifold $\mathcal{M}_{D}$ (defined formally via Incoming and Outgoing Equitable partitions in Sec.˜A.2) represents a significant generalization of the cloning concepts typically discussed in literature. Prior analyses of singularities [Fukumizu and Amari, 2000] or invariant sets [Chen et al., 2023] generally define cloned units by requiring their associated weights to be strictly identical (the block-wise constant condition in our terminology). Our framework proves that invariance under gradient descent holds even under the relaxed condition of equitability, where individual weights may differ as long as specific incoming and outgoing sums are maintained. This significantly broadens the class of structures identified as LoP manifolds.
@@
The cloning LoP manifold naturally lends itself to gradient descent and stochastic gradient descent, regardless of the order which we process the samples, will remain strictly within the manifold. This extends to virtually all variations of gradient descent based optimizations, namely Stochastic Gradient Descent (SGD), SGD with momentum, and Adam, as long as the optimizer is initialized at the onset of cloning. The only exception to this is weigh decay which could break some symmetries. This fact can be empirically observed across our cloning experiments, showing that across a wide range of optimization schemes the model remains trapped onto to the LoP manifold.
Remarkably, the theorem admits a modular version, which allows us to create practical cloning certificate for modern architectures (see Sec.˜A.3 for more details).
@@
To empirically test the validity of the cloning manifold and their potential escape mechanisms we conduct cloning experiments. First, a base model (e.g., an MLP) is trained on a specific task. Subsequently, a larger model is constructed by expanding the base model. This expansion involves increasing the width of the model for MLPs, the number of channels for CNNs and ResNets, and the feature dimension for ViTs. The weights of the cloned model are initialized in such a way that its activations are identical to those of the base model. This effectively creates blocks of units that have identical activations. Next, we train both the base and cloned models on the same task and monitor their training progress through the loss curve, the effective rank of representations, and the cloning $R^{2}$ score. Figure˜2.1 presents the results of such experiments on MLPs, shedding light on the dynamics within and escapes from these LoP manifolds. The empirical validation of these claims, such as demonstrating perfect cloning under specific initializations or the persistence of dead units, can be found in Fig.˜2.1 and Appx.˜B. Notably, despite Adam violating the symmetry conditions required for Theorem˜2.1, the empirical evidence suggests that it frequently fails to escape the manifold. This observation implies the existence of a stronger theory capable of explaining this phenomenon.
@@ 3 Emergence of Loss of Plasticity from Linear–Nonlinear Rank Dynamics
Having established the existence of LoP manifolds, we now discuss the mechanisms within standard training that drive their formation. The optimization process can be seen as a trajectory, beginning with an expansion of representational diversity as features propagate through nonlinear layers and become increasingly decorrelated [Poole et al., 2016]. However, this initial growth is followed by a compression phase, where the network simplifies its representation to retain only the most relevant features for the task. This low-dimensional structure is a key characteristic of neural collapse, where last-layer features for each class converge to their means in a highly organized geometric configuration [Papyan et al., 2020]. This is also consistent with the information bottleneck principle, which describes training as a process of first fitting the data and then compressing the representation to discard irrelevant information [Shwartz-Ziv and Tishby, 2017]. Here, we argue that these two principles derive the model towards the LoP manifolds we identified earlier. Thus, this provides a direct link between these fundamental compression dynamics inherent to deep network optimization, and emergence of LoP.
@@
To diagnose whether features are diversifying or compressing during training, we track a smooth surrogate of the rank of the feature correlation matrix. Exact rank is numerically unstable, because it is not a continuous nor a differentiable map from the matrix space. Therefore, we use differentiable proxies to rank such as Rényi-2 rank, $\mathrm{er}_{2}(M)=(\mathrm{tr}\,M)^{2}/\|M\|_{F}^{2}$, or the Shannon effective rank, $\mathrm{er}(M)=\exp(H(\lambda(M)/\mathrm{tr}\,M))$. Both increase when the eigenmass is evenly distributed or dominated by a few values. Note that both of these surrogates are maximized when matrix $M$ has equal eivenvalues and minimized when it a rank-$1$ matrix. The following theorem offers an insight into how nonlinear layers contribute to the formation of features.
Theorem 3.1 (rank gain across one linear–nonlinear step).
@@
Assume $\phi$ is nonlinear with a Hermite expansion. Let $C$ be the pre-activation correlation matrix with unit diagonal. For an activation $\phi$, define the correlation kernel $K_{\phi}(r)=\mathrm{Corr}(\phi(x),\phi(y))$ where $(x,y)$ are jointly Gaussian with correlation $r$. The nonlinearity acts entrywise on correlations, producing $K_{\phi}(C)$. Its correlation kernel satisfies $K_{\phi}(0)=0$, $K_{\phi}(1)=1$, and $|K_{\phi}(r)|<|r|$ for all $|r|<1$. For any correlation matrix $C$ with unit diagonal the Rényi-2 effective rank obeys
$\frac{\mathrm{er}_{2}(K_{\phi}(C))}{\mathrm{er}_{2}(C)}=\frac{d+\sum_{i\neq j}C_{ij}^{2}}{d+\sum_{i\neq j}K_{\phi}(C_{ij})^{2}}\geq 1,$
and the ratio equals $1$ only when every off-diagonal magnitude $C_{ij}$ is 0 or $1$.
Note that the theorem implies that any correlations $|C_{ij}|\in(0,1)$ strictly increase Rényi-2 rank after the nonlinearity because $K_{\phi}$ is below the identity map everywhere other than those fixed points. Note that this gap also depends on the activation itself, and how nonlinear it is. We can summarize this gap via a scalar that summarizes how strongly the kernel pulls intermediate correlations toward the fixed points $\{0,1\},$ which we term the decorrelation strength $\kappa_{\phi}=\max_{r\in[0,1]}(r^{2}-K_{\phi}(r)^{2})$. Larger $\kappa_{\phi}$ implies a larger rank gain for the same spectrum of $C$. For a formal statement and proof of this statement see Sec.˜A.1.
Implication for emergence of frozen units
So far we implicitly treated pre-activations as standardized. In practice, their first and second moments drift during training. Allowing nonzero means and nonunit variances changes the operating regime of the activation and thus modifies the kernel $K_{\phi}$ and the decorrelation strength $\kappa_{\phi}$. For ReLU, making the effective bias more negative increases nonlinearity and raises $\kappa_{\phi}$. For $\tanh$, increasing effective gain does the same. These are precisely the regimes where the derivative is near zero on most inputs, which explains why training that enhances decorrelation also creates units that are nearly always inactive or saturated. One insight from this connection is that we can unify two empirically observed phenomena: “dead ReLU” and “frozen units” as both symptoms are caused by the same underlying force.
Implication for creation of duplicate or cloned units
Neural collapse is a widely observed endpoint in which the penultimate features are low-rank, and the class means form a simplex or Equiangular Tight Frame (ETF) structure [Papyan et al., 2020]. Under this geometry, the theorem makes a precise prediction. To preserve low rank through the nonlinearity, correlations must lie at kernel fixed points that do not increase effective rank, which means they should be close to 0 or 1. Maintaining rank $C$ therefore requires roughly $C$ orthogonal directions for class structure and dCd-C directions that are near duplicates within classes. Linear layers and the loss encourage this compression and duplication, while the nonlinearity would otherwise expand diversity unless within-class correlations are driven close to 1 and others close to 0.

### v2

@@ Abstract
Deep learning models excel in stationary settings but suffer from loss of plasticity (LoP) in non-stationary environments. While prior literature characterizes LoP through symptoms like rank collapse of representations, it often lacks a mechanistic explanation for why gradient descent fails to recover from these states. This work presents a first-principles investigation grounded in dynamical systems theory, formally defining LoP not merely as a statistical degradation, but as an entrapment of gradient dynamics within invariant sub-manifolds of the parameter space. We identify two primary mechanisms that create these traps: frozen units from activation saturation and cloned-unit manifolds from representational redundancy. Crucially, our framework uncovers a fundamental tension: the very mechanisms that promote generalization in static settings, such as low-rank compression, actively steer the network into these LoP manifolds. We validate our theoretical analysis with numerical simulations and demonstrate how architectural interventions can destabilize these manifolds to restore plasticity.
@@ 1 Introduction
The extraordinary success of back-propagation in training deep neural networks often relies on two implicit assumptions. First, stationarity is assumed: the data distribution encountered during training is similar to the distribution faced during deployment, implying that post-training adaptation is minimal. Second, a single random initialization of network parameters is the main source of diversity and exploration potential—a resource that is progressively consumed by optimization and rarely replenished. These assumptions falter when an artificial agent must operate and learn continuously within an environment characterized by changing dynamics or evolving task distributions. This scenario, commonly referred to as continual or lifelong learning, presents the stability-plasticity dilemma [Abraham and Robins, 2005, Chaudhry et al., 2018], demanding a system stable enough to retain acquired knowledge yet plastic enough to integrate new information.
Empirically, standard deep networks subjected to long sequences of tasks or drifting data streams exhibit a sharp decline in learning capability [Dohare et al., 2024, Berariu et al., 2021, Nikishin et al., 2022]. This phenomenon, termed Loss of Plasticity (LoP), is distinct from catastrophic forgetting: while forgetting erases old knowledge, LoP prevents the acquisition of new knowledge. Common symptoms are well-documented: exploding weight magnitudes [Nikishin et al., 2022], the emergence of “dead” units [Sokar et al., 2023, Lyle et al., 2022], and a collapse in the effective rank of representations [Papyan et al., 2020, Huh et al., 2023, Kumar et al., 2020]. While recent works have linked LoP to back-propagation mechanics [Dohare et al., 2024] or NTK rank degradation [Lyle et al., 2023], these observations are largely descriptive. They characterize the symptoms of the failing network but stop short of a formal mechanistic explanation for the persistence of the non-plastic state.
A fundamental question remains: Why exactly does gradient descent fail to recover from these states? If LoP is merely a bad configuration, why do gradients not steer the model back toward useful regions? For example, while many works have identified low rank as a hallmark of LoP, why is the network unable to recover feature diversity under a new task distribution? Our work revisits the dynamics of gradient descent through the lens of dynamical systems theory to answer this. We posit that LoP is not simply a degradation of statistics, like rank or weight norm, but a topological entrapment under gradient descent. We seek to identify the specific geometric structures underlying the optimization landscape that act as “sinks” for gradient trajectories, restricting the model from accessing its full degrees of freedom.
The central contribution of this work is the formalization of LoP manifolds: invariant sub-manifolds of the parameter space that act as “traps” for gradient descent (Fig.˜1.1). Once the optimization process steers parameters into these regions, the gradient flow becomes strictly tangent to the manifold, rendering escape impossible without external intervention. This formulation fills a critical theoretical gap: while prior literature has extensively cataloged the symptoms of plasticity loss (e.g., weight explosion, rank collapse), our framework identifies the mathematical and geometric principles that provably make these states persistent and irreversible. Furthermore, we show the driver towards this entrapment—a fundamental rank-plasticity tension. We show that the very mechanisms that promote generalization in static settings, specifically, the compression of features toward low-rank structures, are the primary forces that steer the network into these LoP manifolds. Thus, the dynamics that maximize performance on the current task inadvertently construct the barrier to adaptability for future tasks.
[…]
Figure 1.1: Conceptual illustration of LoP as a topological trap. The high-dimensional parameter space $\Theta$ contains low-dimensional, invariant LoP manifolds $\mathcal{M}\subset\Theta$ . Left (Independent tasks): During standard learning, trajectories starting at $\theta_{0}$ can explore the full space (e.g., $\theta_{B},\theta_{C}$), though implicit biases may eventually drive them toward simpler representations on $\mathcal{M}$ ($\theta_{A}$). Right (Sequential tasks): In continual learning, early tasks drive parameters onto the manifold ($\theta_{0}\to\theta_{A}$). Once trapped, gradients for subsequent tasks become strictly tangent to $\mathcal{M}$, constraining future dynamics to this restricted subspace ($\theta_{A}\to\theta_{B}\to\theta_{C}\to\dots$) and preventing the recovery of plasticity needed for new data distributions.
Summary of Contributions.
In this work, we formalize the mechanics of plasticity loss and validate our theory across varied architectures. Our specific contributions are:
  • A Dynamical Systems Definition of LoP (Sec.˜2). We move beyond symptomatic definitions to propose a formal definition of LoP as the entrapment of optimization trajectories within specific invariant sub-manifolds of the parameter space.
  • Identification of Trap Mechanisms (Sec.˜2). We theoretically characterize two primary classes of invariant manifolds: Frozen-Unit Manifolds ($\mathcal{M}_{F}$), caused by activation saturation, and Cloned-Unit Manifolds ($\mathcal{M}_{C}$), caused by representational redundancy. We prove that standard gradient-based optimization cannot escape these manifolds once entered.
  • The Rank-Plasticity Tension (Sec.˜3). We derive a theoretical connection between feature rank dynamics and plasticity. We show that properties widely considered beneficial for generalization (e.g., neural collapse) act as attractive forces toward LoP manifolds.
  • Empirical Validation and Mitigation (Sec.˜4). We validate our framework through extensive experiments on MLPs, ResNets, and ViTs. We further analyze how architectural interventions (normalization) and targeted perturbations (noise injection, dropout) affect the stability of these manifolds, providing a principled basis for recovering plasticity.
@@ 2 LoP Manifolds: Traps for Gradient Descent
In this section, we lay the groundwork for our analysis by defining Loss of Plasticity (LoP) within a dynamical systems framework and recalling the standard definitions for feed-forward neural networks and back-propagation. The stability of LoP manifolds, a crucial concept for understanding their persistence, will be discussed later in Sec.˜2. Let $\theta\in\Theta\subseteq\mathbb{R}^{p}$ represent the parameters of a neural network. We consider training on a stream of data $\{(x_{i},y_{i})\}_{i=1}^{N}$ using gradient descent or its stochastic variants. The objective is typically to minimize a loss $\sum_{i=1}^{N}\mathcal{L}(\hat{y}_{\theta}(x_{i}),y_{i}),$ which we can succinctly refer to as $\mathcal{L}(\theta)$. This allows us to define LoP based on the trajectory of parameters $\theta(t)$ in the parameter space $\Theta$ as driven by the negative gradient in the loss landscape. As illustrated in the right panel of Fig.˜1.1, we are interested in scenarios where this trajectory becomes confined to a lower-dimensional subspace $\mathcal{M}\subset\Theta$.
@@
Remark 2.1.
Definition 2.1 is stated in terms of continuous-time gradient flow $\dot{\theta}=-\nabla_{\theta}\mathcal{L}(\theta)$: the tangency condition $\nabla_{\theta}\mathcal{L}(\theta)\in T_{\theta}\mathcal{M}$ ensures that the flow starting in $\mathcal{M}$ remains in $\mathcal{M}$. For a general curved manifold this need not imply invariance under discrete updates $\theta_{k+1}=\theta_{k}-\eta_{k}\nabla_{\theta}\mathcal{L}(\theta_{k})$, as discretization step can lead to an escape from the manifold. However, for affine LoP manifolds, both gradient flow and (stochastic) gradient descent dynamics are invariant once they enter $\mathcal{M}$, as the discretization step cannot may lead to an escape.
@@
Note that both LoP manifolds $\mathcal{M}_{F}$ and $\mathcal{M}_{C}$ are defined as linear LoP manifolds in the sense of Definition˜2.1. Formal proofs and further details are provided in Sec.˜A.3.
@@
It is important to note that the Duplicate Manifold $\mathcal{M}_{D}$ (defined formally via Incoming and Outgoing Equitable partitions in Sec.˜A.3) represents a significant generalization of the cloning concepts typically discussed in literature. Prior analyses of singularities [Fukumizu and Amari, 2000] or invariant sets [Chen et al., 2023] generally define cloned units by requiring their associated weights to be strictly identical (the block-wise constant condition in our terminology). Our framework proves that invariance under gradient descent holds even under the relaxed condition of equitability, where individual weights may differ as long as specific incoming and outgoing sums are maintained. This significantly broadens the class of structures identified as LoP manifolds.
@@
The cloning LoP manifold naturally lends itself to gradient descent and stochastic gradient descent, regardless of the order which we process the samples, will remain strictly within the manifold. This extends to virtually all variations of gradient descent based optimizations, namely Stochastic Gradient Descent (SGD), SGD with momentum, and Adam, as long as the optimizer is initialized at the onset of cloning. The only exception to this is weight decay which could break some symmetries. This fact can be empirically observed across our cloning experiments, showing that across a wide range of optimization schemes the model remains trapped onto to the LoP manifold.
Remarkably, the theorem admits a modular version, which allows us to create practical cloning certificate for modern architectures (see Sec.˜A.4 for more details).
@@
To empirically test the validity of the cloning manifold and their potential escape mechanisms we conduct cloning experiments. First, a base model (e.g., an MLP) is trained on a specific task. Subsequently, a larger model is constructed by expanding the base model. This expansion involves increasing the width of the model for MLPs, the number of channels for CNNs and ResNets, and the feature dimension for ViTs. The weights of the cloned model are initialized in such a way that its activations are identical to those of the base model. This effectively creates blocks of units that have identical activations. Next, we train both the base and cloned models on the same task and monitor their training progress through the loss curve, the effective rank of representations, and the cloning $R^{2}$ score. Figure˜2.1 presents the results of such experiments on MLPs, shedding light on the dynamics within and escapes from these LoP manifolds. The empirical validation of these claims, such as demonstrating perfect cloning under specific initializations or the persistence of dead units, can be found in Fig.˜2.1 and Appx.˜B. Notably, Remark˜2.4 predicts that Adam, despite fundamental differences with SGD, also fails to escape the manifold, which is validated by our experiments.
@@ 3 Emergence of Loss of Plasticity from Linear–Nonlinear Rank Dynamics
Having established the existence of LoP manifolds, we now discuss the mechanisms within standard training that drive their formation. The optimization process naturally follows a trajectory beginning with an expansion of representational diversity, as features propagate through nonlinear layers and become increasingly decorrelated [Poole et al., 2016]. However, this expansion is counterbalanced by a drive toward geometric simplicity in the final phase of training, as captured by neural collapse [Papyan et al., 2020], where within-class variability vanishes and features converge to a low-rank subspace spanned by class means. We argue that this geometric collapse—while beneficial for generalization in static settings—inherently steers the model toward the low-rank LoP manifolds we identified. Thus, the very dynamics that maximize separability on the current task provide the direct mechanism for the emergence of LoP.
@@
To diagnose whether features are diversifying or compressing during training, we track a smooth surrogate of the rank of the feature correlation matrix. Exact rank is numerically unstable, because it is not a continuous nor a differentiable map from the matrix space. Therefore, we use differentiable proxies to rank such as Rényi-2 rank, $\mathrm{er}_{2}(M)=(\mathrm{tr}\,M)^{2}/\|M\|_{F}^{2}$, or the Shannon effective rank, $\mathrm{er}(M)=\exp(H(\lambda(M)/\mathrm{tr}\,M))$. Both increase when the eigenmass is evenly distributed and decrease when dominated by a few values. Note that both of these surrogates are maximized when matrix $M$ has equal eivenvalues and minimized when it a rank-$1$ matrix. The following theorem offers an insight into how nonlinear layers contribute to the formation of features.
Theorem 3.1 (Rank gain and decorrelation potential).
@@
Assume $\phi\in L^{2}(\mathcal{N}(0,1))$ is nonlinear and normalized such that $\mathbb{E}[\phi(Z)]=0$ and $\mathrm{Var}(\phi(Z))=1$. Let $C$ be a pre-activation correlation matrix. For an activation $\phi$, define the correlation kernel $K_{\phi}(r)=\mathrm{Corr}(\phi(x),\phi(y))$ where $(x,y)$ are jointly Gaussian with correlation $r$. The nonlinearity acts entrywise on correlations, producing $K_{\phi}(C)$. We define the decorrelation strength $\gamma_{\phi}=\frac{1-K_{\phi}^{\prime}(0)}{1+K_{\phi}^{\prime}(0)}$ and the decorrelation potential $\Psi(C)=\sum_{i\neq j}C_{ij}^{2}(1-C_{ij}^{2})$. The Rényi-2 effective rank is non-decreasing according to the bound below:
$\frac{\mathrm{er}_{2}(K_{\phi}(C))}{\mathrm{er}_{2}(C)}\;\geq\;1+\gamma_{\phi}\frac{\Psi(C)}{\|C\|_{F}^{2}}.$
The ratio is strictly larger than $1$ when the potential $\Psi(C)$ is vanishing (all off-diagonal $C_{ij}\in\{0,\pm 1\}$) and the the activation is linear ($\gamma_{\phi}$ is zero).
Note that the theorem implies that any correlations within range $|C_{ij}|\in(0,1)$ provide a potential $\Psi(C)>0$ that the nonlinearity consumes to increase rank. The rate of this increase is governed by the scalar $\gamma_{\phi}\in[0,1]$, depends on the activation and quantifies the nonlinearity’s ability to de-correlate inputs and thereby increase rank. Larger $\gamma_{\phi}$ guarantees a larger rank gain for the same input spectrum. For a rigorous derivation, see Sec.˜A.1.
Implication for emergence of frozen units.
In practice, the first and second moments of pre-activations drift during training, modulating the effective activation function. This modulation directly impacts the decorrelation strength $\gamma_{\phi}$. We find that $\gamma_{\phi}$ is monotonic with respect to the saturation parameters: for ReLU, making the effective bias more negative increases $\gamma_{\phi}$; for $\tanh$, increasing the effective gain does the same. Consequently, optimizing for feature diversity (high decorrelation strength) creates a pressure to push units into regimes where the derivative vanishes on most inputs. This explains why training that effectively restores rank also creates “frozen” or “dead” units.
Implication for creation of duplicate or cloned units.
Neural collapse is a widely observed endpoint in which the penultimate features are low-rank, and the class means form a simplex or Equiangular Tight Frame (ETF) structure [Papyan et al., 2020]. The theorem makes a precise prediction regarding this state: for the rank to stabilize (ratio $\approx 1$) while retaining nonlinearity ($\gamma_{\phi}>0$), the decorrelation potential $\Psi(C)$ must vanish. This forces correlations to converge to the fixed points 0 (orthogonal features) or $\pm 1$ (aligned or anti-aligned features). Thus, the dynamics of rank expansion naturally drive the network toward a mixture of orthogonal subspaces and cloned units, matching the geometry of LoP manifolds.
@@
Remark 3.1.
Our analysis explicitly links LoP to the “Rich” or feature learning regime. In the “Lazy” learning regime (associated with the Neural Tangent Kernel), weights remain close to initialization, maintaining high-rank connectivity and avoiding feature compression. Conversely, the “Rich” regime is characterized by significant weight evolution and the emergence of low-rank structures [Chizat et al., 2019, Woodworth et al., 2020]. The frozen units and cloned manifolds we identify are essentially symptoms of this feature learning process carried to an extreme. Thus, LoP can be viewed as a pathology specific to the Rich regime: the same mechanism that enables deep networks to learn efficient representations (data compression) eventually traps them in low-rank manifolds that hinder future adaptability.